The three dimensions of a rectangular box are height, length, and width. The us of the dimensions is 18 feet. The height equals four fifths the sum of the length and the width. It also equals four times the difference of the length and the width. What are the dimensions of the box?
Answer by mananth(16946)

sum of dimensions = 18 feet
l+w+h=18
(l+w)=18-h
.
.
second condition
h= 4/5(l+w)
replace l+w by (18-h)
h=4/5(18-h)
multiply equation by 5
5h= 4(18-h)
5h=72-4h
5h+4h=72
9h=72
h=8
h=4(l-w)
8=l-w
l=8+w
l+w+h=18
8+w+w+8=18
2w+16=18
2w=18-16
2w=2
w=1
so Length =8+w= 9 feet (w=1)
the dimensions are 9 by 8 by 1 feet